# Automated Privacy Information Annotation in Large Language Model Interactions

### v1

@@ Appendix I Discussion of Main Results
"<|QUERY|>"
Prompts for in-context learning (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
{CJK}UTF8gbsn 隐私短语应当有具体的指代,而不是一个泛指的词汇。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是一些例子:
{CJK}UTF8gbsn <|CASE|>例子结束。
{CJK}UTF8gbsn 用户当前的请求是:
"<|QUERY|>"
Table 24: Prompt template for in-context learning of privacy information summarization task in English and Chinese dataset, where <|INPUT|> is the input of user query and <|CASE|>is structured examples from training set.
Prompts for in-context learning (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
@@
"<|QUERY|>"
Prompts for in-context learning (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
{CJK}UTF8gbsn 然后,对于每个提取的短语,总结该短语泄露的具体隐私信息
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是一些例子:
{CJK}UTF8gbsn <|CASE|>例子结束。
{CJK}UTF8gbsn 用户当前的请求是:
"<|QUERY|>"
Table 25: Instructions for training models for privacy leakage classification task, where <|INPUT|> is the input of user query.
Instruction for English datasets.
[…]
System:
Judge whether the query that reveal any privacy-sensitive information of the user, or user related people or things.
@@
Query: "<|INPUT|>"
Instruction for Chinese datasets.
[…]
System:
{CJK}UTF8gbsn 判断Query是否泄漏了用户、用户相关人员或事物的隐私敏感信息
@@
{CJK}UTF8gbsn Query: "<|INPUT|>"
Table 26: Instructions for training models for privacy phrase extraction task, where <|INPUT|> is the input of user query.
Instruction for English datasets.
[…]
System:
@@
Query: "<|INPUT|>"
Instruction for Chinese datasets.
[…]
System:
{CJK}UTF8gbsn 从Query中提取出能够揭示用户、用户相关人员或事物的隐私敏感信息的短语
@@
{CJK}UTF8gbsn Query: "<|INPUT|>"
Table 27: Instructions for training models for privacy information summarization task, where <|INPUT|> is the input of user query.
Instruction for English datasets.
[…]
System:
@@
Query: "<|INPUT|>"
Instruction for Chinese datasets.
[…]
System:
{CJK}UTF8gbsn 从Query中提取出能够揭示用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 然后,针对每个提取的短语,总结该短语泄露的具体隐私信息
@@
Query: "<|INPUT|>"

### v2

@@ Appendix F Ethical Statement
”¡—QUERY—¿”
Prompts for in-context learning (Chinese). 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
隐私短语应当有具体的指代,而不是一个泛指的词汇。
请一步一步思考(Think step-by-step)。
给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
以下是一些例子:
¡—CASE—¿例子结束。
用户当前的请求是:
”¡—QUERY—¿”
Figure 22. Prompt template for in-context learning of privacy phrase extraction task in English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—CASE—¿is structured examples from the training set.
Prompts for in-context learning (English). Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
@@
”¡—QUERY—¿”
Prompts for in-context learning (Chinese). 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
然后,对于每个提取的短语,总结该短语泄露的具体隐私信息
请一步一步思考(Think step-by-step)。
给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
以下是一些例子:
¡—CASE—¿例子结束。
用户当前的请求是:
”¡—QUERY—¿”
Figure 23. Prompt template for in-context learning of privacy information summarization task in English and Chinese dataset, where ¡—INPUT—¿ is the input of the user query and ¡—CASE—¿is structured examples from the training set.
Instruction for English datasets. System:
Judge whether the query that reveals any privacy-sensitive information of the user, or user related people or things.
@@
Query: ”¡—INPUT—¿”
Instruction for Chinese datasets. System:
判断Query是否泄漏了用户、用户相关人员或事物的隐私敏感信息
@@
Query: ”¡—INPUT—¿”
Figure 24. Instructions for training models for privacy leakage classification task, where ¡—INPUT—¿ is the input of the user query.
Instruction for English datasets. System:
@@
Query: ”¡—INPUT—¿”
Instruction for Chinese datasets. System:
从Query中提取出能够揭示用户、用户相关人员或事物的隐私敏感信息的短语
@@
Query: ”¡—INPUT—¿”
Figure 25. Instructions for training models for privacy phrase extraction task, where ¡—INPUT—¿ is the input of user query.
Instruction for English datasets. System:
@@
Query: ”¡—INPUT—¿”
Instruction for Chinese datasets. System:
从Query中提取出能够揭示用户、用户相关人员或事物的隐私敏感信息的短语。
然后,针对每个提取的短语,总结该短语泄露的具体隐私信息
@@
Query: ”¡—INPUT—¿”
Figure 26. Instructions for training models for privacy information summarization task, where ¡—INPUT—¿ is the input of the user query.